# Monte Carlo sampling with integrator snippets

### v1

@@ Abstract
Assume interest is in sampling from a probability distribution $\mu$ defined on $(\mathsf{Z},\mathscr{Z})$. We develop a framework to construct sampling algorithms taking full advantage of numerical integrators of ODEs, say $\psi\colon\mathsf{Z}\rightarrow\mathsf{Z}$ for one integration step, to explore $\mu$ efficiently and robustly. The popular Hybrid/Hamiltonian Monte Carlo (HMC) algorithm [17, 29] and its derivatives are example of such a use of numerical integrators. However we show how the potential of integrators can be exploited beyond current ideas and HMC sampling in order to take into account aspects of the geometry of the target distribution. A key idea is the notion of integrator snippet, a fragment of the orbit of an ODE numerical integrator $\psi$, and its associate probability distribution $\bar{\mu}$, which takes the form of a mixture of distributions derived from $\mu$ and $\psi$. Exploiting properties of mixtures we show how samples from $\bar{\mu}$ can be used to estimate expectations with respect to $\mu$. We focus here primarily on Sequential Monte Carlo (SMC) algorithms, but the approach can be used in the context of Markov chain Monte Carlo algorithms as discussed at the end of the manuscript. We illustrate performance of these new algorithms through numerical experimentation and provide preliminary theoretical results supporting observed performance.
@@ 1 Overview and motivation: SMC sampler with HMC
Assume interest is in sampling from a probability distribution $\mu$ on a probability space $(\mathsf{Z},\mathscr{Z})$. The main ideas of sequential Monte Carlo (SMC) samplers to simulate from $\mu$ are (a) to define a sequence of probability distributions $\{\mu_{n},n\in\llbracket 0,P\rrbracket\}$ on $(\mathsf{Z},\mathscr{Z})$ where $\mu_{P}=\mu$, $\mu_{0}$ is chosen by the user, simple to sample from and the sequence $\{\mu_{n},n\in\llbracket P-1\rrbracket\}$ “interpolates” $\mu_{0}$ and $\mu_{P}$, (b) and to propagate a cloud of samples $\{z_{n}^{(i)}\in\mathsf{Z},i\in\llbracket N\rrbracket\}$ for $n\in\llbracket 0,P\rrbracket$ to represent $\{\mu_{n},n\in\llbracket 0,P\rrbracket\}$ using an importance sampling/resampling mechanism [16]. Notation and definitions used throughout this paper can be found in Appendix A.
@@
We adopt this measure theoretic formulation of importance weights out of necessity in order to take into account situations such as when $\mu_{n-1}$ and $\mu_{n}$ both have densities with respect to the Lebesgue measure but $M_{n}$ is a Metropolis-Hastings (MH) update, which does not possess such a density – more details are provided in Appendices A-B but can be omitted to understand how the algorithms considered in the manuscript proceed.
@@
Given $\{M_{n},n\in\llbracket P\rrbracket\}$, theoretically optimal choice of $\{L_{n-1},n\in\llbracket P\rrbracket\}$ is well understood but tractability is typically obtained by assuming that $M_{n}$ is $\mu_{n-1}$-invariant, or considering approximations of $\{L_{n-1},n\in\llbracket P\rrbracket\}$ and that $M_{n}$ is $\mu_{n}$-invariant. This makes Markov chain Monte Carlo (MCMC) kernels very attractive choices for $M_{n}$, and the use of measure theoretic tools inevitable.
@@ 12            Set $z_{n}^{(i)}=\tilde{z}_{n}^{(a_{i})}$
A possible choice of MCMC kernel is that of the hybrid Monte Carlo method, a MH update using a discretization of Hamilton’s equations [17, 29], which can be thought of as a particular instance of a more general strategy relying on numerical integrators of ODEs possessing properties of interest. More specifically, assume that interest is in sampling $\pi$ defined on $(\mathsf{X},\mathscr{X})$. First the problem is embedded into that of sampling from the joint distribution $\mu({\rm d}z):=\pi\otimes\varpi\big{(}{\rm d}z\big{)}=\pi({\rm d}x)\varpi({\rm d}v)$ defined on $(\mathsf{Z},\mathscr{Z})=(\mathsf{X}\times\mathsf{V},\mathscr{X}\otimes\mathscr{V})$, where $v$ is an auxiliary variable facilitating sampling. Following the SMC framework we set $\mu_{n}({\rm d}z):=\pi_{n}\otimes\varpi_{n}\big{(}{\rm d}z\big{)}$ for $n\in\llbracket 0,P\rrbracket,$ a sequence of distributions on $(\mathsf{Z},\mathscr{Z})$ with $\pi_{P}=\pi$, $\{\pi_{n},n\in\llbracket 0,P-1\rrbracket\}$ probabilities on $\big{(}\mathsf{X},\mathscr{X}\big{)}$ and $\{\varpi_{n},n\in\llbracket 0,P\rrbracket\}$ on $\big{(}\mathsf{V},\mathscr{V}\big{)}$. With $\psi\colon\mathsf{Z}\rightarrow\mathsf{Z}$ an integrator of an ODE of interest, one can use $\psi^{k}(z)$ for some $k\in\mathbb{N}$ as a proposal in a MH update mechanism; $v$ is a source of randomness allowing “exploration”, resampled every now and then. Again, hereafter we let $z=:(x,v)\in\mathsf{X}\times\mathsf{V}$ be the corresponding components of $z$.
@@
and possess the important property that $H_{n}(x_{t},v_{t})=H_{n}(x_{0},v_{0})$ for $t\geq 0$. The corresponding leapfrog integrator is given, for some $\varepsilon>0$, by
@@
${}_{\textsc{b}}\psi(x,v):=\big{(}x,v-\tfrac{1}{2}\varepsilon\nabla U_{n}(x)\big{)},\quad{}_{\textsc{a}}\psi_{n}(x,v)=(x+\varepsilon\,v,v)\,.$
We point out that, with the exception of the first step, only one evaluation of $\nabla U(x)$ is required per integration step since the rightmost ${}_{\textsc{b}}\psi$ in (3) recycles the last computation from the last iteration. Let $\sigma\colon\mathsf{Z}\rightarrow\mathsf{Z}$ such that for any $f\colon\mathsf{Z}\rightarrow\mathsf{Z}$, $f\circ\sigma(x,v)=f(x,-v)$, it is standard to check that $\psi_{n}^{-1}=\sigma\circ\psi_{n}\circ\sigma$ and note that $\mu_{n}\circ\sigma=\mu_{n}$ in this particular case. In its most basic form the hybrid Monte Carlo MH update leaving $\mu_{n}$ invariant proceeds as follows, for $(z,A)\in\mathsf{Z}\times\mathscr{Z}$
@@
In turn these quantities offer the promise of being able to exploit points used to generate the snippet while preserving accuracy of the estimators of interest, through importance sampling or reweighting. For example one may consider virtual HMC updates i.e. given $(z,A)\in\mathsf{Z}\times\mathscr{Z}$ and $k\in\llbracket P-1\rrbracket$ define
@@
Noting that $\mu_{n}P_{n,k}=\mu_{n}$ one deduces that for $z\sim\mu_{n}$
$\frac{1}{T+1}\sum_{k=0}^{T}\alpha_{n}(z;k)f\circ\psi_{n}^{k}(z)+\bar{\alpha}_{n}(z;k)f(z)\,,$
is an unbiased estimator of $\mu(f)$. This post-processing procedure is akin to waste-recycling [14] and its generalizations but, crucially, does not affect the dynamic of the Markov chain generating the samples. An alternative algorithm exploiting the snippet fully, with an active effect on the dynamics, could use the following mixture of Markov chain transition kernels [23, 29, 22],
which is also related to the strategy adopted in [11] with the extra chance algorithm. As we shall see our work shares the same objective but we adopt a strategy more closely related to [28]; see Section 5 for a more detailed discussion.
The present paper is concerned with exploring such recycling procedures in the context of SMC samplers, although the ideas we develop can be straightforwardly applied to particle filters in the context of state-space models and to MCMC as discussed later. The manuscript is organised as follows.
In Section 2 we first provide a high level description of particular instances of the class of algorithms considered and provide a justification through reinterpretation as standard SMC algorithms in Subsection 2.2. In Subsection 2.3 we discuss direct extensions of our algorithms, some of which we explore in the present manuscript. This work has links with related recent attempts in the literature [33, 14, 38]; these links are discussed and contrasted with our work in Subsection 2.4 where some of the motivations behind these algorithms are also discussed. Initial exploratory simulations demonstrating the interest of our approach are provided in Subsections 2.5 and 2.6.
In Section 3 we introduce the more general framework of Markov snippets Monte Carlo and associated formal justifications. Subsection 3.3 details the link with the scenario considered in Section 2. In Subsection 3.4 we provide general results facilitating the practical calculation of some of the Radon-Nikodym involved, highlighting why some of the usual constraints on mutation and backward kernels in SMC can be lifted here.
In Section 4 we provide elements of a theoretical analysis explaining expected properties of the algorithms proposed in this manuscript, although a fully rigorous theoretical analysis is beyond the present methodological contribution.
In Section 5 we explore the use of some of the ideas developed here in the context of MCMC algorithms and establish links with earlier suggestions, such as “windows of states” techniques proposed in the context of HMC [28, 29]. Notation, definitions and basic mathematical background can be found in Appendices A-B.
@@ 2 An introductory example
Assume interest is in sampling from a probability distribution $\mu$ on $(\mathsf{Z},\mathscr{Z})$ as described above Example 1 using an SMC sampler relying on the leapfrog integrator of Hamilton’s equations. As in the previous section we introduce an interpolating sequence of distributions $\{\mu_{n},n\in\llbracket 0,P\rrbracket\}$ on $(\mathsf{Z},\mathscr{Z})\}$ and assume for now the existence of densities for $\{\mu_{n},n\in\llbracket 0,P\rrbracket\}$ with respect to a common measure $\upsilon$, say the Lebesgue measure on $\mathbb{R}^{2d}$, denoted $\mu_{n}(z):={\rm d}\mu_{n}/{\rm d}\upsilon(z)$ for $z\in\mathsf{Z}$ and $n\in\llbracket 0,P\rrbracket$ and denote $\psi_{n}$ the corresponding integrator, which again is measure preserving in this setup.
@@ 17            Set $\bar{z}_{n}^{(j)}:=(\bar{x}_{n-1}^{(j)},\bar{v}_{n-1}^{(j)})=z_{n-1,a_{j}}^{(b_{j})}$
The SMC sampler-like algorithm in Alg. 2 therefore involves propagating $N$ “seed” particles $\{z_{n-1}^{(i)},i\in\llbracket N\rrbracket\}$, with a mutation mechanism consisting of the generation of $N$ integrator snippets $\mathsf{z}:=\big{(}z,\psi_{n}(z),\psi_{n}^{2}(z),\ldots,\psi_{n}^{T}(z)\big{)}$ started at every seed particle $z\in\{z_{n-1}^{(i)},i\in\llbracket N\rrbracket\}$, resulting in $N\times(T+1)$ particles which are then whittled down to a set of $N$ seed particles using a standard resampling scheme; after rejuvenation of velocities this yields the next generation of seed particles–this is illustrated in Fig. 1. This algorithm should be contrasted with standard implementations of SMC samplers where, after resampling, a seed particle normally gives rise to a single particle in the mutation step, in Fig.  1 the last state on the snippet. Intuitively validity of the algorithm follows from the fact that if $\big{\{}(z_{n-1}^{(i)},1),i\in\llbracket N\rrbracket\big{\}}$ represent $\mu_{n-1}$, then $\big{\{}\big{(}z_{n-1,k}^{(i)},\bar{w}_{n,k}(z_{n-1}^{(i)})\big{)}),(i,k)\in\llbracket N\rrbracket\times\llbracket 0,T\rrbracket\big{\}}$ represents $\mu_{n}$ in the sense that for $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ summable, one can use the approximation
@@
where the self-normalization of the weights is only required in situations where the ratio $\mu_{n}(z)/\mu_{n-1}(z)$ is only known up to a constant. We provide justification for the correctness of Alg. 2 and the estimator (7) by recasting the procedure as a standard SMC sampler targetting a particular sequence of distributions in Section 2.2 and using properties of mixtures. Direct generalizations are provided in Section 2.3.
@@
2.2 Justification outline

### v2

@@ Abstract
Assume interest is in sampling from a probability distribution $\mu$ defined on $(\mathsf{Z},\mathscr{Z})$. We develop a framework for sampling algorithms which takes full advantage of ODE numerical integrators, say $\psi\colon\mathsf{Z}\rightarrow\mathsf{Z}$ for one integration step, to explore $\mu$ efficiently and robustly. The popular Hybrid Monte Carlo (HMC) algorithm [29, 48] and its derivatives are examples of such a use of numerical integrators. A key idea developed here is that of sampling integrator snippets, that is fragments of the orbit of an ODE numerical integrator $\psi$, and the definition of an associated probability distribution $\bar{\mu}$ such that expectations with respect to $\mu$ can be estimated from integrator snippets distributed according to $\bar{\mu}$. The integrator snippet target distribution $\bar{\mu}$ takes the form of a mixture of pushforward distributions which suggests numerous generalisations beyond mappings arising from numerical integrators, e.g. normalising flows. Very importantly this structure also suggests new principled and robust strategies to tune the parameters of integrators, such as the discretisation stepsize, effective integration time, or number of integration steps, in a Leapfrog integrator.
We focus here primarily on Sequential Monte Carlo (SMC) algorithms, but the approach can be used in the context of Markov chain Monte Carlo algorithms. We illustrate performance and, in particular, robustness through numerical experiments and provide preliminary theoretical results supporting observed performance.
@@ 1 Overview and motivation: SMC sampler with HMC
Assume interest is in sampling from a probability distribution $\mu$ on a probability space $(\mathsf{Z},\mathscr{Z})$. The main ideas of sequential Monte Carlo (SMC) samplers to simulate from $\mu$ are (a) to define a sequence of probability distributions $\{\mu_{n},n\in\llbracket 0,P\rrbracket\}$ on $(\mathsf{Z},\mathscr{Z})$ where $\mu_{P}=\mu$, $\mu_{0}$ is chosen by the user, simple to sample from and the sequence $\{\mu_{n},n\in\llbracket P-1\rrbracket\}$ “interpolates” $\mu_{0}$ and $\mu_{P}$, (b) and to propagate a cloud of samples $\{z_{n}^{(i)}\in\mathsf{Z},i\in\llbracket N\rrbracket\}$ for $n\in\llbracket 0,P\rrbracket$ to represent $\{\mu_{n},n\in\llbracket 0,P\rrbracket\}$ using an importance sampling/resampling mechanism [28].
@@
More details are provided in Appendices A-B concerning the existence and definition of these weights, but can be omitted on a first reading.
@@
Given $\{M_{n},n\in\llbracket P\rrbracket\}$, theoretically optimal choice of $\{L_{n-1},n\in\llbracket P\rrbracket\}$ is well understood but tractability is typically obtained by assuming that $M_{n}$ is $\mu_{n-1}$-invariant, or considering approximations of $\{L_{n-1},n\in\llbracket P\rrbracket\}$ and that $M_{n}$ is $\mu_{n}$-invariant. This makes Markov chain Monte Carlo (MCMC) kernels very attractive choices for $M_{n}$.
@@ 12            Set $z_{n}^{(i)}=\tilde{z}_{n}^{(a_{i})}$
A possible choice of MCMC kernel is the Hybrid Monte Carlo (HMC) method, a Metropolis-Hastings (MH) update using a discretisation of Hamilton’s equations [29, 48] to update the state, a particular instance of the use of numerical integrators of ODEs in this context. More specifically, assume that interest is in sampling $\pi$ defined on $(\mathsf{X},\mathscr{X})$. First the problem is embedded into that of sampling from the joint distribution $\mu({\rm d}z):=\pi\otimes\varpi\big{(}{\rm d}z\big{)}=\pi({\rm d}x)\varpi({\rm d}v)$ defined on $(\mathsf{Z},\mathscr{Z})=(\mathsf{X}\times\mathsf{V},\mathscr{X}\otimes\mathscr{V})$, where $v$ is an auxiliary variable facilitating sampling. Following the SMC framework we set $\mu_{n}({\rm d}z):=\pi_{n}\otimes\varpi_{n}\big{(}{\rm d}z\big{)}$ for $n\in\llbracket 0,P\rrbracket,$ a sequence of distributions on $(\mathsf{Z},\mathscr{Z})$ with $\pi_{P}=\pi$, $\{\pi_{n},n\in\llbracket 0,P-1\rrbracket\}$ probabilities on $\big{(}\mathsf{X},\mathscr{X}\big{)}$ and $\{\varpi_{n},n\in\llbracket 0,P\rrbracket\}$ on $\big{(}\mathsf{V},\mathscr{V}\big{)}$. With $\psi\colon\mathsf{Z}\rightarrow\mathsf{Z}$ an integrator of an ODE of interest, one can use $\psi^{k}(z)$ for some $k\in\mathbb{N}$ as a proposal in a MH update mechanism; $v$ is a source of randomness allowing “exploration”, resampled every now and then. Again, hereafter we let $z=:(x,v)\in\mathsf{X}\times\mathsf{V}$ be the corresponding components of $z$.
@@
and possess the important property that $H_{n}(x_{t},v_{t})=H_{n}(x_{0},v_{0})$ for $t\geq 0$. The corresponding leapfrog integrator is given, for some $\epsilon>0$, by
@@
${}_{\textsc{b}}\psi(x,v):=\big{(}x,v-\tfrac{1}{2}\epsilon\nabla U_{n}(x)\big{)},\quad{}_{\textsc{a}}\psi_{n}(x,v)=(x+\epsilon\,v,v)\,.$
We point out that, with the exception of the first step, only one evaluation of $\nabla U(x)$ is required per integration step since the rightmost ${}_{\textsc{b}}\psi$ in (3) recycles the last computation from the last iteration. Let $\sigma\colon\mathsf{Z}\rightarrow\mathsf{Z}$ be such that for any $f\colon\mathsf{Z}\rightarrow\mathsf{Z}$, $f\circ\sigma(x,v)=f(x,-v)$, then in its most basic form the HMC update leaving $\mu_{n}$ invariant proceeds as follows, for $(z,A)\in\mathsf{Z}\times\mathscr{Z}$
@@
In turn these quantities offer the promise of being able to exploit points used to generate the snippet while preserving accuracy of the estimators of interest, through importance sampling or reweighting. For example an algorithm exploiting the snippet fully could use the following mixture of $\mu_{n}$-invariant Markov chain transition kernels [42, 48, 39],
where for $(z,A)\in\mathsf{Z}\times\mathscr{Z}$ and $k\in\llbracket P-1\rrbracket$
@@
As we shall see our work shares the same objective but we adopt a strategy more closely related to [47] (see Appendix I for a more detailed discussion) which however leads to a comparison between states along a snippet. Further, an advantage of SMC samplers is that they provide population wide information about tuning parameters, in particular of numerical integrators, therefore suggesting robust self-tuning procedures.
The manuscript is organised as follows. In Section 2 we first provide a high level description of particular instances of the class of algorithms considered and provide a justification through reinterpretation as standard SMC algorithms in Subsection 2.2. In Subsection 2.4 we discuss direct extensions of our algorithms, some of which we explore in the present manuscript. This work has links to related recent attempts in the literature [55, 26, 61]; these links are discussed and contrasted with our work in Subsection 2.5 where some of the motivations behind these algorithms are also discussed. Early exploratory simulations demonstrating the interest of our approach are provided in Appendices E-F. Our main focus is however on the development of adaptive algorithms in Section 3 where, focussing on the Leapfrog integrator, we develop novel criteria and updating strategies to adapt the discretisation stepsize and integration time by taking advantage of the population-wide information available; in simulation the resulting algorithms display remarkable robustness on the examples we have considered We draw some conclusions in Section 4.
The manuscript contains a number of appendices, containing supporting material and additional developments. Notation, definitions and basic mathematical background can be found in Appendices A-B. Proofs for Section 2 are provided in Appendix C. In Appendix D we introduce the more general framework of Markov snippets Monte Carlo and associated formal justifications; the algorithms used in Appendix E.2 rely on some of these extensions. Appendix D.3 details the link with the scenario considered in Section 2. In Appendix D.4 we provide general results facilitating the practical calculation of some of the Radon-Nikodym involved, highlighting why some of the usual constraints on mutation and backward kernels in SMC can be lifted here. Appendix G contains ancillary supporting details for Section 3. In Appendix H we provide elements of a theoretical analysis explaining expected properties of the algorithms proposed in this manuscript, although a fully rigorous theoretical analysis is beyond the present methodological contribution. In Appendix I we explore the use of some of the ideas developed here in the context of MCMC algorithms and establish links with earlier suggestions, such as “windows of states” techniques proposed in the context of HMC [47, 48].
@@ 2 An introductory example
We still aim to sample from a probability distribution $\mu$ on $(\mathsf{Z},\mathscr{Z})$ as described above Example 1 using an SMC sampler relying on the leapfrog integrator of Hamilton’s equations. As in the previous section we introduce an interpolating sequence of distributions $\{\mu_{n},n\in\llbracket 0,P\rrbracket\}$ on $(\mathsf{Z},\mathscr{Z})\}$ and assume for now the existence of densities for $\{\mu_{n},n\in\llbracket 0,P\rrbracket\}$ with respect to a common measure $\upsilon$, say the Lebesgue measure on $\mathbb{R}^{2d}$, denoted $\mu_{n}(z):={\rm d}\mu_{n}/{\rm d}\upsilon(z)$ for $z\in\mathsf{Z}$ and $n\in\llbracket 0,P\rrbracket$ and denote $\psi_{n}$ the corresponding integrator, which again is measure preserving in this setup.
@@ 17            Set $\bar{z}_{n}^{(j)}:=(\bar{x}_{n-1}^{(j)},\bar{v}_{n-1}^{(j)})=z_{n-1,a_{j}}^{(b_{j})}$
The SMC sampler-like algorithm in Alg. 2 therefore involves propagating $N$ “seed” particles $\{z_{n-1}^{(i)},i\in\llbracket N\rrbracket\}$, with a mutation mechanism consisting of the generation of $N$ integrator snippets $\mathsf{z}:=\big{(}z,\psi_{n}(z),\psi_{n}^{2}(z),\ldots,\psi_{n}^{T}(z)\big{)}$ started at every seed particle $z\in\{z_{n-1}^{(i)},i\in\llbracket N\rrbracket\}$, resulting in $N\times(T+1)$ particles which are then whittled down to a set of $N$ seed particles using a standard resampling scheme; after rejuvenation of velocities this yields the next generation of seed particles–this is illustrated in Fig. 1; a more realistic depiction of Alg. 2 in the context of Example 1 with $\pi$ a two dimensional distribution is presented in Fig. 2. This algorithm should be contrasted with standard implementations of SMC samplers where, after resampling, a seed particle normally gives rise to a single particle in the mutation step, in Fig.  1 the last or first state on the snippet when using an MH as mutation kernel. Intuitively validity of the algorithm follows from the fact that if $\big{\{}(z_{n-1}^{(i)},1),i\in\llbracket N\rrbracket\big{\}}$ represent $\mu_{n-1}$, then $\big{\{}\big{(}z_{n-1,k}^{(i)},\bar{w}_{n,k}(z_{n-1}^{(i)})\big{)}),(i,k)\in\llbracket N\rrbracket\times\llbracket 0,T\rrbracket\big{\}}$ represents $\mu_{n}$ in the sense that for $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ summable, one can use the approximation
@@
where the self-normalization of the weights is only required in situations where the ratio $\mu_{n}(z)/\mu_{n-1}(z)$ is only known up to a constant.
Simple numerical experiments demonstrating robustness and efficiency of our approach are provided in Appendices E-F. However our primary objective, beyond establishing correctness, is to show that our proposed framework naturally lends itself to novel adaptation strategies, for example of the stepsize $\epsilon$ and integration time $T$; see Section 3. To that purpose we first provide justification for the correctness of Alg. 2 and the estimator (7) by recasting the procedure as a standard SMC sampler targetting a particular sequence of distributions in Section 2.2 and using properties of mixtures. Direct generalizations are provided in Section 2.4 and the adaptation strategies of Section 3 then follow.
@@
Figure 2: Evolution of the $x$ components of $5$ seed particles and their integrator snippets ($T=15$) targetting a normal distribution $\pi$ defined on $\mathbb{R}^{2}$, using the leapfrog integrator of Example 1.
2.2 Justification